OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

What methods of solving systems of linear equations have you been taught/ do you want to use on this one?

OpenStudy (anonymous):

Ok, look at the second equation: 4x + y = 57 Whenever you have a variable already alone (ie it has no number beside it) that usually makes life easier for you. In this case, try to isolate y (in other words, put it on one side of the equation, and all the rest on the other side) Go ahead :D

OpenStudy (anonymous):

4x + y = 57 Now add (-4x) to both sides 4x + y - 4x = 57 - 4x 4x - 4x + y = 57 - 4x y = 57 - 4x Get me so far? :)

OpenStudy (anonymous):

x - 2y = 21 4x + y = 57 We take the second equation, and multiply everything by 2 2(4x + y) = 2(57) 8x + 2y = 114 We replace the second equation with this one 7x - 2y = 21 8x + 2y = 114 Now we add up all those on the left side, and all those on the right 7x - 2y + 8x + 2y = 21 + 114 15x = 135 (the y got cancelled out!) x = 9 (divided both sides by 15) So now that we know that x = 9, then we substitute it for the x in the first equation 7x - 2y = 21 7(9) - 2y = 21 63 - 2y = 21 - 2y = 21 - 63 - 2y = -42 y = 21 (dividing both sides by -2) So x=9 and y=21 To be safe, let's try it with the second equation as well 4x + y = 57 4(9) + 21 = 57 36 + 21 = 57 57 = 57 (it works! ) Thus, we're sure that x = 9 and y = 21 END Hope that helps And remember to study the process I went through to get to the answer